A local hamburger shop sold a combined total of 434 hamburgers and cheeseburgers on Thursday. There were 66 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Thursday?
I have tried subtracting 66 from 434 and its not a choice on the multiple question, not sure how to solve this problem, thanks for your help :-)

Answer by ReadingBoosters(3246)

x = hamburgers
y = cheeseburgers
x+y=434
66 fewer cheeseburgers than hamburgers
y = x - 66
Substitute y into the first equation
x + (x-66) = 434
2x = 434 + 66
2x = 500
x = 250 hamburgers sold